solve.
f(x) = (x+3)^2 find x so that f(x) = 10

Answer by stanbon(75887)

solve.
f(x) = (x+3)^2 find x so that f(x) = 10
--------------------------------------------
(x+3)^2 = 10
---
(x+3)^2-10 = 0
Factor the difference of squares:
[x+3+sqrt(10)][x+3-sqrt(10)] = 0
x = -3-sqrt(10)) or x = -3+sqrt(10)
=======================================
Cheers,
Stan H.